If two buildings are 155 meters and 248 meters tall, respectively, what is the scale factor of the smaller building to the larger building?
Answer is 5:8 but how did one arrive at that answer??????

Answer by jim_thompson5910(35256)

155/248 = (5*31)/(8*31) = 5/8

So the scale factor is 5:8
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------